# BTL-UI: Blink-Think-Link Reasoning Model for GUI Agent

### v2

@@ Abstract
In the field of AI-driven human-GUI interaction automation, while rapid advances in multimodal large language models and reinforcement fine-tuning techniques have yielded remarkable progress, a fundamental challenge persists: their interaction logic significantly deviates from natural human-GUI communication patterns. To fill this gap, we propose “Blink-Think-Link” (BTL), a brain-inspired framework for human-GUI interaction that mimics the human cognitive process between users and graphical interfaces. The system decomposes interactions into three biologically plausible phases: (1) Blink - rapid detection and attention to relevant screen areas, analogous to saccadic eye movements; (2) Think - higher-level reasoning and decision-making, mirroring cognitive planning; and (3) Link - generation of executable commands for precise motor control, emulating human action selection mechanisms. Additionally, we introduce two key technical innovations for BTL framework: (1) Blink Data Generation - an automated annotation pipeline specifically optimized for blink data, and (2) BTL Reward – the first rule-based reward mechanism that enables reinforcement learning driven by both process and outcome. Building upon this framework, we develop a GUI agent model named BTL-UI, which demonstrates consistent state-of-the-art performance across both static GUI understanding and dynamic interaction tasks in comprehensive benchmarks. These results provide conclusive empirical validation of the framework’s efficacy in developing advanced GUI Agents.
@@ 1 Introduction
The automation of graphical user interface (GUI) interactions constitutes a pivotal milestone in developing genuinely intelligent digital assistants wang2024gui ; hu2024agents ; nguyen2024gui . Recent breakthroughs in large vision-language models (VLMs) wang2024qwen2 ; bai2025qwen2 and reinforcement learning fine-tuning techniques have substantially enhanced agents’ capabilities in natural language command interpretation, visual element perception, and multi-step task execution through human-like reasoning hong2024cogagent ; cheng2024seeclick .
However, current mainstream systems primarily adopt two approaches. The first relies on supervised fine-tuning (SFT) to align model behavior with task objectives, but this method faces two major limitations: a strong dependence on large-scale expert-labeled data and limited generalization capability when faced with out-of-distribution scenarios. The second approach involves rule-based reinforcement fine-tuning (RFT) guo2025deepseek , as shown in Figure 1 (a), which enhances generalization in complex tasks by using a structured reasoning format: intermediate cognitive steps are encapsulated within <think> tags, and final decisions are expressed through <answer> tags. While effective in improving task performance, these methods qin2025ui ; xia2025gui ; liu2025infigui exhibit two critical shortcomings: (1) significant deviation from natural human-GUI interaction patterns, and (2) excessive focus on interaction outcomes while lacking effective process-oriented reward mechanisms.
Figure 1: Framework comparison of previous “Think-Answer” and “Blink-Think-Link” in GUI tasks for RFT. Specifically, colorful text is supervised by rule-based reinforcement learning. And different colors of text indicate different reward rules. The previous “Think-Answer” framework is optimized by format reward, action type reward, and corresponding args reward. And our “Blink-Think-Link” framework is optimized by dual format reward, blink reward, and link reward.
@@
Overall, the main contributions are summarized as follows:
@@
    We propose BTL (Blink-Think-Link), an innovative framework that simulates the human cognitive process in human-GUI interaction by explicitly modeling how users perceive, process, and act upon interface elements.
@@
    We develop BTL-UI, a GUI agent trained via BTL framework, extensive experiment results demonstrate that the model achieves state-of-the-art performance across multiple GUI benchmarks.
Figure 2: Overall framework of BTL. We adopt GRPO to optimize the proposed BTL. Firstly, the base model generates NN completions for a given GUI task sample. Furthermore, GRPO computes the relative advantages within a group of completions, eliminating the need for manually annotated data. Finally, the policy model updates parameters under the guidance of relative advantages and the KL divergence constraint.
@@ 2.1 GUI Agents
Autonomous agents powered by LLMs and VLMs have recently garnered considerable scholarly interest due to their interactive functionalities. For GUI tasks, earlier systems relied on LLMs to read and interpret structured representations such as HTML and accessibility trees. However, since icons, images, diagrams, and their spatial relationships are difficult to express in such structured languages, LLM-based agents often perform poorly hong2024cogagent ; nong2024mobileflow ; song2024visiontasker . Therefore, VLM-based agents are introduced to perceive visual GUI signals directly with better performance hu2024agents ; liu2024autoglm ; shen2024falcon ; tang2025think ; christianos2024lightweight ; zheng2025vem . For instance, UGround gou2024navigating developed a specialized GUI grounding model for GUI element localization. OS-Atlas wu2024atlas proposed a foundational model for GUI agents by interpreting human intentions and predicting actions in the form of function calls. Aguvis xu2024aguvis integrated explicit planning and reasoning within the model, enhancing its ability to navigate and interact with complex digital environments autonomously. UI-TARS qin2025ui combines GUI-related pretraining with task-level reasoning fine-tuning to better capture the complexity of GUI interactions. Although research on VLM-based GUI agents has made impressive progress, they mainly follow the training paradigm of SFT, which directly imitates the ground-truth actions provided in the curated data.
@@ 2.2 Reinforcement Fine-Tuning
With the advent of rule-based reinforcement learning approaches such as OpenAI-o1 jaech2024openai and DeepSeek-R1 guo2025deepseek , recent studies have demonstrated that RFT enhances the model’s reasoning capabilities and affords it greater generalization ability liu2025visual . Subsequent approaches liu2025visual ; zhan2025vision ; tan2025reason have introduced this paradigm to VLMs. For instance, Vision-R1 zhan2025vision combined a vision criterion-driven reward function and a progressive rule refinement strategy to enhance VLM’s object localization capabilities. Visual-RFT liu2025visual adopted the GRPO-based reinforcement learning strategy to enhance the visual perception and grounding ability of VLMs. VLM-R1 shen2025vlm demonstrated that RFT with small amounts of high-quality data can enable VLMs to solve complex vision-language tasks.
For GUI tasks, UI-R1 lu2025ui and GUI-R1 xia2025gui introduced rule-based reinforcement learning frameworks that require minimal expert supervision, demonstrating competitive performance. InfiGUI-R1 liu2025infigui further advanced the field by bridging reactive execution and deliberative reasoning via the Actor2Reasoner architecture. However, existing RFT-based GUI agents predominantly adopt rule-based reinforcement learning, which focus on final outcomes and lack intermediate process guidance, often overlooking key aspects of human cognition and interaction.
@@ 3 Method
In this section, we introduce BTL, a new framework grounded in cognitive science theory, with its core concept derived from the “Blink-Think-Link” paradigm observed in human-GUI interactions. The framework is shown in Figure 2. We explore the implementation details of this framework through the following components: Preliminaries, Blink Data Generation, BTL Reward, and Policy Optimization.
@@ 3.1 Preliminaries
The interaction between a GUI agent and its environment can be naturally formulated as a Markov Decision Process (MDP), defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{Z},\mathcal{T},\mathcal{O}\rangle$. Here, $\mathcal{S}$ denotes the state space representing possible screen states; $\mathcal{A}$ is the action space, encompassing interaction types such as clicking, typing, and scrolling; $\mathcal{Z}$ is the observation space, including screenshots or structured UI representations; $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ defines the probability of transitioning from one state to another given an action; and $\mathcal{O}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{Z}$ ispecifies the probability of receiving a particular observation given a state and an action.
During task execution, at each discrete time step $t$, the agent receives an input tuple $(z_{t},u,h)$, where $z_{t}\in\mathcal{Z}$ is the current screen state, $u$ refers to the global task instruction, and $h$ is its interaction history. The BTL process can then be formalized as a structured policy function $\mathcal{F}$:
@@
where $o_{t}$ denotes the BTL output at time $t$, consisting of: $b_{t}$—visual attention regions, $d_{t}$—vreasoning and decision trace, $a_{t}$—vthe final action to be executed. Each action $a_{t}=(\alpha_{t},\delta_{t})\in\mathcal{A}$ is composed of an action type $\alpha_{t}$ (e.g., click) and its corresponding parameters $\delta_{t}$ (e.g., coordinates, text input). Upon executing $a_{t}$, the environment transitions to a new state $z_{t+1}$, and the process repeats until the task is completed or a terminal condition is met.
@@ 3.2 Blink Data Generation
One of the core innovations of the BTL framework is its ability to simulate the human mechanism of rapidly locating ROIs in a visual scene during the blink phase. To achieve this, we propose an automated Blink data generation pipeline that annotates ROIs on the screenshot corresponding to the user instruction in the MDP. As illustrated in Figure 3, the pipeline consists of two main stages. A parsing model lu2024omniparserpurevisionbased first processes the raw screenshot to extract semantic UI elements. Then, an analysis model bai2025qwen2 is used to evaluate the visual salience and contextual relevance of these elements, enabling filtering and prioritization to generate the final ROI annotations.
Specifically, in the first stage, we extract individual UI elements such as buttons, icons, and text fields, annotating each with bounding box coordinates, type, and semantic captions. These annotations form a structured representation of the screen state, enabling bottom-up human-like interpretation of the GUI. The output of this stage is a comprehensive list of elements, denoted as $\mathcal{E}=\{e_{1},e_{2},\dots,e_{n}\}$, where each $e_{k}=\{{id},{bbox},{type},{caption},{interactivity}\}$ represents the attributes related to the element. This foundational representation serves as the input for subsequent filtering and prioritization steps that model instruction-directed visual attention.
In the second stage, we employ Qwen2.5-VL 32B bai2025qwen2 to simulate top-down attention by filtering and ranking elements based on visual saliency and task relevance. Guided by the task instruction $u$ and the interaction history $h$, the model dynamically selects a subset of elements from the parsed list $\mathcal{E}$. This selection process can be formulated as:
@@ 3.3 BTL Reward
Effective GUI agents must excel at both interface grounding and long-horizon planning. To this end, we design a three-component, rule‑based BTL reward $R_{\text{BTL}}$ scheme that mirrors the human Blink–Think–Link cognitive cycle:
@@
Dual Format Reward. Following previous works guo2025deepseek ; liu2025visual leveraging format rewards to encourage predefined templates for easy answer extraction, we introduce a dual format reward to evaluate whether the generated output adheres to both the expected structural template and content. Specifically, the template check function $f_{\text{template}}$ is used to check whether the generated completions meet the “Blink-Think-Link” three-stage grammatical structure. Furthermore, the content check function $f_{\text{content}}$ is adopted to evaluate whether the blink content complies with the XML format and the link content complies with the JSON format, which facilitates the parsing of trajectory planning and actions with corresponding arguments. We adopt a binary reward scheme, assigning a reward of 1 only when the prediction $o_{i}$ fully satisfies both format and content criteria as follows:
@@
Blink Reward. This component incentivizes rapid, accurate localization of interface elements relevant to the instruction $u$. From the agent’s prediction $o_{i}$, we extract a set of ROIs $P_{i}=\{p_{i}^{x}\}$ and compare them to ground-truth annotations $G_{i}=\{g_{i}^{x}\}$ (see §3.2). Using the Hungarian matcher kuhn1955hungarian $M(\cdot,\cdot,\tau)$ with IoU threshold $\tau$, we define the matched index set:
@@
Link Reward. The link phase assesses the agent’s ability to generate a fully coherent, executable command. Recent RFT-based GUI agents xia2025gui ; liu2025infigui ; lu2025ui always split the reward of the predicted action into a reward for the action type and a reward for the action args (e.g, click coordinates or input text). However, this kind of reward will split an action into two independent contents, which is not in line with human cognition. At the same time, this staged action reward will cause reward hacking, which prevents the agent’s understanding of the designed action space. Thus, we employ a strict binary criterion: the agent receives a reward only if both the action type and its associated arguments are exactly correct. Formally, the link reward is defined as:
@@ 3.4 Advantage Computation and Parameter Update
As shown in Figure 2, we adopt GRPO to optimize the proposed BTL. Since its supervision is based solely on the final outcome, GRPO is especially suited for tasks with explicit, objective answers. Furthermore, GRPO significantly reduces memory overhead for VLMs by removing the reward models or value models in other performance optimization methods schulman2017proximal .
@@
$A_{i}=\frac{R_{i}-mean(\{R_{j}\}_{j=1}^{N})}{std(\{R_{j}\}_{j=1}^{N})}.$ (8)
@@
$\mathcal{J}_{GRPO}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{\pi_{\theta}(o_{i}|z_{i},u,h)}{\pi_{\theta_{old}}(o_{i}|z_{i},u,h)}A_{i}-\beta\mathcal{KL}(\pi_{\theta}(o_{i}|z_{i},u,h)|\pi_{ref}(o_{i}|z_{i},u,h))\right),$ (9)

### v3

@@ Abstract
In the field of AI-driven human-GUI interaction automation, while rapid advances in multimodal large language models and reinforcement fine-tuning techniques have yielded remarkable progress, a fundamental challenge persists: their interaction logic significantly deviates from natural human-GUI communication patterns. To address this gap, we propose Blink–Think–Link (BTL), a brain-inspired framework for human-GUI interaction that mimics the human cognitive process between users and graphical interfaces. The system decomposes interactions into three biologically plausible phases: (1) Blink - rapid detection and attention to relevant screen areas, analogous to saccadic eye movements; (2) Think - higher-level reasoning and decision-making, mirroring cognitive planning; and (3) Link - generation of executable commands for precise motor control, emulating human action selection mechanisms. Additionally, we introduce two key technical innovations for BTL framework: (1) Blink Data Generation - an automated annotation pipeline specifically optimized for blink data, and (2) BTL Reward – the first rule-based reward mechanism that enables reinforcement learning driven by both process and outcome. Building upon this framework, we develop a GUI agent model named BTL-UI, which demonstrates competitive performance across both static GUI understanding and dynamic interaction tasks in comprehensive benchmarks. These results provide conclusive empirical validation of the framework’s efficacy in developing advanced GUI agents.
@@ 1 Introduction
Automation of graphical user interface (GUI) interactions constitutes a pivotal milestone in developing genuinely intelligent digital assistants wang2024gui ; hu2024agents ; nguyen2024gui . Recent breakthroughs in large vision-language models (VLMs) wang2024qwen2 ; bai2025qwen2 and reinforcement learning fine-tuning techniques have substantially improved agents’ capabilities in natural language command interpretation, visual element perception, and multi-step task execution through human-like reasoning hong2024cogagent ; cheng2024seeclick .
However, current mainstream systems adopt mainly two approaches. The first relies on supervised fine-tuning (SFT) to align model behavior with task objectives, but this method faces two major limitations: a strong dependence on large-scale expert-labeled data and limited generalization capability when faced with out-of-distribution scenarios. The second approach involves rule-based reinforcement fine-tuning (RFT) guo2025deepseek , as shown in Figure 1 (a), which enhances generalization in complex tasks by using a structured reasoning format: intermediate cognitive steps are encapsulated within <think> tags, and final decisions are expressed through <answer> tags. Although effective in improving task performance, these methods qin2025ui ; xia2025gui ; liu2025infigui exhibit two critical shortcomings: (1) significant deviation from natural human-GUI interaction patterns, and (2) excessive focus on interaction outcomes while lacking effective process-oriented reward mechanisms.
Figure 1: Framework comparison of previous Think-Answer and Blink-Think-Link in GUI tasks for RFT. Specifically, colorful text is supervised by rule-based reinforcement learning. And different colors of text indicate different reward rules. The previous “Think-Answer” framework is optimized by format reward, action type reward, and corresponding args reward. And our Blink-Think-Link framework is optimized by dual format reward, blink reward, and link reward.
@@
In general, the main contributions are summarized as follows:
@@
    We propose BTL (Blink-Think-Link), an innovative framework that simulates the human cognitive process in the human-GUI interaction by explicitly modeling how users perceive, process, and act upon interface elements.
@@
    We develop BTL-UI, a GUI agent trained via the BTL framework, and extensive experiment results demonstrate that the model achieves competitive performance across multiple GUI benchmarks.
Figure 2: Overall framework of BTL. We adopt Group Relative Policy Optimization (GRPO) to optimize the proposed BTL. Firstly, the base model generates NN completions for a given GUI task sample. Furthermore, GRPO computes the relative advantages within a group of completions, eliminating the need for manually annotated data. Finally, the policy model updates parameters under the guidance of relative advantages and the KL divergence constraint.
@@ 2.1 GUI Agents
Autonomous agents powered by large language models (LLMs) and VLMs have recently garnered considerable scholarly interest due to their interactive functionalities. For GUI tasks, earlier systems relied on LLMs to read and interpret structured representations such as HTML and accessibility trees. However, since icons, images, diagrams, and their spatial relationships are difficult to express in such structured languages, agents based on LLM often perform poorly hong2024cogagent ; nong2024mobileflow ; song2024visiontasker . Therefore, VLM-based agents have been introduced to perceive visual GUI signals directly with better performance hu2024agents ; liu2024autoglm ; shen2024falcon ; tang2025think ; christianos2024lightweight ; zheng2025vem . For example, UGround gou2024navigating developed a specialized GUI grounding model for GUI element localization. OS-Atlas wu2024atlas proposed a foundational model for GUI agents by interpreting human intentions and predicting actions in the form of function calls. Aguvis xu2024aguvis integrated explicit planning and reasoning within the model, enhancing its ability to navigate and interact with complex digital environments autonomously. UI-TARS qin2025ui combines GUI-related pretraining with task-level reasoning fine-tuning to better capture the complexity of GUI interactions. Although research on VLM-based GUI agents has made impressive progress, they mainly follow the SFT training paradigm, which directly mimics the ground-truth actions provided in the curated data.
@@ 2.2 Reinforcement Fine-Tuning
With the advent of rule-based reinforcement learning approaches such as OpenAI-o1 jaech2024openai and DeepSeek-R1 guo2025deepseek , recent studies have demonstrated that RFT improves the reasoning abilities of the model and provides greater generalizability liu2025visual . Subsequent approaches liu2025visual ; zhan2025vision ; tan2025reason have introduced this paradigm to VLMs. For example, Vision-R1 zhan2025vision combined a vision criterion-driven reward function and a progressive rule refinement strategy to enhance VLM’s object localization capabilities. Visual-RFT liu2025visual adopted the reinforcement learning strategy to enhance visual perception and grounding ability of VLMs. VLM-R1 shen2025vlm demonstrated that RFT with small amounts of high-quality data can enable VLMs to solve complex vision-language tasks.
For GUI tasks, UI-R1 lu2025ui and GUI-R1 xia2025gui introduced rule-based reinforcement learning frameworks that require minimal expert supervision, demonstrating competitive performance. InfiGUI-R1 liu2025infigui further advanced the field by bridging reactive execution and deliberative reasoning through the Actor2Reasoner architecture. However, existing RFT-based GUI agents predominantly adopt rule-based reinforcement learning, which focus on final outcomes and lack intermediate process guidance, often overlooking key aspects of human cognition and interaction.
@@ 3 Method
In this section, we introduce BTL, a new framework grounded in cognitive science theory, with its core concept derived from the Blink-Think-Link paradigm observed in human-GUI interactions. The framework is shown in Figure 2. We detail the implementation details of this framework through the following components: Preliminaries, Blink Data Generation, BTL Reward, and Policy Optimization.
@@ 3.1 Preliminaries
The interaction between a GUI agent and its environment can be naturally formulated as a Markov Decision Process (MDP), defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{Z},\mathcal{T},\mathcal{O}\rangle$. Here, $\mathcal{S}$ denotes the state space representing possible screen states; $\mathcal{A}$ is the action space that encompasses interaction types such as clicking, typing, and scrolling; $\mathcal{Z}$ is the observation space, including screenshots or structured UI representations; $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ defines the probability of transitioning from one state to another given an action; and $\mathcal{O}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{Z}$ specifies the probability of receiving a particular observation given a state and an action.
During task execution, at each discrete time step $t$, the agent receives an input tuple $(z_{t},u,h)$, where $z_{t}\in\mathcal{Z}$ is the current state of the screen, $u$ refers to the global task instruction and $h$ is its interaction history. The BTL process can then be formalized as a structured policy function $\mathcal{F}$:
@@
where $o_{t}$ denotes the BTL output at time $t$, consisting of: $b_{t}$—visual attention regions, $d_{t}$—reasoning and decision trace, $a_{t}$—the final action to be executed. Each action $a_{t}=(\alpha_{t},\delta_{t})\in\mathcal{A}$ is composed of an action type $\alpha_{t}$ (e.g., click) and its corresponding parameters $\delta_{t}$ (e.g., coordinates, text input). Upon execution of $a_{t}$, the environment transitions to a new state $z_{t+1}$, and the process repeats until the task is completed or a terminal condition is met.
@@ 3.2 Blink Data Generation
One of the core innovations of the BTL framework is its ability to simulate the human mechanism of rapidly locating ROIs in a visual scene during the blink phase. To achieve this, we propose an automated Blink data generation pipeline that annotates ROIs on the screenshot corresponding to the user instruction in the MDP. As illustrated in Figure 3, the pipeline consists of two main stages. A parsing model lu2024omniparserpurevisionbased first processes the raw screenshot to extract semantic UI elements. Then, an analysis model bai2025qwen2 is used to evaluate the visual importance and contextual relevance of these elements, allowing filtering and prioritization to produce the final ROI annotations.
Specifically, in the first stage, we extract individual UI elements such as buttons, icons, and text fields, annotating each with bounding box coordinates, type, and semantic captions. These annotations form a structured representation of the state of the screen, enabling bottom-up human-like interpretation of the GUI. The output of this stage is a comprehensive list of elements, denoted $\mathcal{E}=\{e_{1},e_{2},\dots,e_{n}\}$, where each $e_{k}=\{\text{id}_{k},\text{bbox}_{k},\text{type}_{k},\text{caption}_{k},\text{interactivity}_{k}\}$ represents the attributes related to the element. This foundational representation serves as the input for subsequent filtering and prioritization steps that model instruction-directed visual attention.
In the second stage, we employ Qwen2.5-VL-32B bai2025qwen2 to simulate top-down attention by filtering and ranking elements based on visual saliency and task relevance. Oriented by task instruction $u$ and interaction history $h$, the model dynamically selects a subset of elements from the filtered list $\mathcal{E}$. This selection process can be formulated as:
@@ 3.3 BTL Reward
Effective GUI agents must excel at both interface grounding and long-horizon planning. To this end, we design a three-component rule-based reward scheme, denoted as $R_{\text{BTL}}$ that mirrors the human Blink–Think–Link cognitive cycle:
@@
Dual Format Reward. Following previous work guo2025deepseek ; liu2025visual that leverages format rewards to encourage predefined templates for easy answer extraction, we introduce a dual format reward to evaluate whether the generated output adheres to both the expected structural template and content. Specifically, the template check function $f_{\text{template}}$ is used to check whether the generated completions meet the Blink-Think-Link three-stage grammatical structure. Furthermore, the content check function $f_{\text{content}}$ is adopted to evaluate whether the blink content complies with the XML format and the link content complies with the JSON format, which facilitates the parsing of trajectory planning and actions with corresponding arguments. We adopt a binary reward scheme, assigning a reward of 1 only when the prediction $o_{i}$ fully satisfies both format and content criteria as follows:
@@
Blink Reward. This component incentivizes the rapid and accurate localization of the interface elements relevant to the instruction $u$. From the agent’s prediction $o_{i}$, we extract a set of ROIs $P_{i}=\{p_{i}^{x}\}$ and compare them to ground-truth annotations $G_{i}=\{g_{i}^{x}\}$ (see §3.2). We adopt the Hungarian matcher kuhn1955hungarian $M(\cdot,\cdot,\tau)$, a classical assignment algorithm used to compute the optimal one-to-one matching between predicted and ground-truth bounding boxes based on IoU scores, under a given threshold $\tau$.
@@
Link Reward. The link phase assesses the agent’s ability to generate a fully coherent executable command. Recent RFT-based GUI agents xia2025gui ; liu2025infigui ; lu2025ui always split the reward of the predicted action into a reward for the action type and a reward for the action args (e.g. click coordinates or input text). However, this kind of reward will split an action into two independent contents, which is not in line with human cognition. At the same time, this staged action reward will cause reward hacking, which prevents the agent from understanding the designed action space. Thus, we employ a strict binary criterion: the agent receives a reward only if both the action type and its associated arguments are exactly correct. Formally, the link reward is defined as:
@@ 3.4 Advantage Computation and Parameter Update
As shown in Figure 2, we adopt Group Relative Policy Optimization (GRPO) to optimize the proposed BTL. Since its supervision is based solely on the final outcome, GRPO is particularly suited for tasks with explicit, objective answers. Furthermore, GRPO significantly reduces memory overhead for VLMs by removing the reward models or value models in other performance optimization methods schulman2017proximal .
@@
$A_{i}=\frac{R_{i}-\text{mean}(\{R_{j}\}_{j=1}^{N})}{\text{std}(\{R_{j}\}_{j=1}^{N})}.$ (8)
@@
$\mathcal{J}_{GRPO}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{\pi_{\theta}(o_{i}|z_{i},u,h)}{\pi_{\theta_{\text{old}}}(o_{i}|z_{i},u,h)}A_{i}-\beta\cdot\mathcal{KL}(\pi_{\theta}(o_{i}|z_{i},u,h)|\pi_{\text{ref}}(o_{i}|z_{i},u,h))\right),$ (9)